OpenStudy (theghostlycreeper):

What is the value of v? 29v+4−23v=5

KjSaif (kjsaif):

Step 1: Simplify both sides of the equation. 29v+4−23v=5 Simplify: (Show steps) 6v+4=5 Step 2: Subtract 4 from both sides. 6v+4−4=5−4 6v=1 Step 3: Divide both sides by 6. 6v/6=1/6

KjSaif (kjsaif):

v=1/6

KjSaif (kjsaif):

:)

OpenStudy (theghostlycreeper):

wait 1/6

KjSaif (kjsaif):

yes